# Robust Optimal Safe and Stability Guaranteeing Reinforcement Learning Control for Quadcopter

### v1

@@ Abstract
Recent advances in deep learning has provided new data-driven ways of controller design, to replace the traditional manual synthesis and certification approaches. Employing neural network (NN) as controllers however presents its own challenge: that of certifying stability, due to their inherent complex nonlinearity, and while NN controllers have demonstrated high performance in complex systems, they often lack formal stability guarantees. This issue is further accentuated for critical nonlinear applications such as of unmanned aerial vehicles (UAVs), complicating their stability guarantees, whereas a lack of stability assurance raises the risk of critical damage or even complete failure under a loss of control. In this study, we improve a Robust, Optimal, Safe and Stability Guaranteed Training (ROSS-GT) method of [1] to design an NN controller for a quadcopter flight control. The approach ensures closed-loop system stability by finding a Lyapunov function, and providing a safe initial state domain that remains invariant under the control and guarantees stability to an equilibrium within it. Stability guaranteeing constraints are derived from the sector bound of the system nonlinearity and of its parameters and disturbance variations, in the form of a Lipschitz bound for a NN control. The control performance is further optimized by searching over the class of stability-guaranteeing controllers to minimize the reference tracking error and the control costs.
@@ I INTRODUCTION
In recent decades, neural networks (NNs) based data-driven approach has achieved great progress, finding applications across diverse fields. One promising usecase is in the design of NN controllers for dynamical systems, where a neural network maps states or observations to actions. Unlike conventional feedback controllers, which rely either on predefined structures with domain-specific expertise for their tuning or control Lyapunov functions requiring manual exploration, NN controllers are trained through data-driven approaches. For instance, supervised imitation learning uses images from a human operator and trains a neural network to imitate those actions [2],[3]. Similarly, in reinforcement learning (RL), policy networks are optimized to maximize the expected cumulative rewards from control sequences [4].
@@
While NN controllers have demonstrated good performance in many applications, their deployment in safety-critical systems, such as unmanned aerial vehicles (UAVs), remains limited due to the lack of formal stability guarantees. Yet another important consideration for controllers is robustness—the ability of a system to withstand disturbances, sensor noise, or model uncertainties without loss of stability. In real-world environments characterized by various uncertainties, robust control is essential to ensure reliable operation, which again requires a formal certification that in general is lacking for NN controllers.
@@
This paper enhances and applies the Robust, Optimal, Safe and Stability-Guaranteed Training (ROSS-GT) method of [1] to design and train an NN controller for a quadcopter, a widely used form of UAV. The ROSS-GT method consists of two stages: First, it extracts the constraints to ensure robust stability and safety for a closed-loop system as well as a safe initial state domain that is invariant and guarantees converges to an equilibrium within, based on the sector bounds of system nonlinearities and bounds for parametric variations and disturbances, in form of a Lipschitz bound for the neural network (i.e., their output needs to bounded by the input magnitude and the Lipschitz bound gain). The maximal permissible robust, safe, and stabilizing Lipschitz bound and the initial safe state domain ensuring asymptotic stability under bounded parametric uncertainty and disturbance are derived by iteratively solving a linear matrix inequality (LMI) test. Second, it finds an optimal NN controller satisfying the Lipschitz bound obtained in the first step to minimize control cost and reference trajectory deviation error using reinforcement learning.
@@
The original ROSS-GT method [1] employs the infinity norm to derive the Lipschitz bound for the NN controller. However, this approach often yields overly conservative estimates for the NN Lipschitz constant, particularly for NN with many neurons, making it impractical. In this paper, we propose an enhancement, replacing the infinity norm with the Euclidean norm, demonstrating that it still satisfies the LMI constraints, while reducing the above mentioned conservativeness. Furthermore, we utilize a semidefinite programming-based method for the estimation of Lipschitz constant of a neural network [5] that again enhances the method in [1], enabling less conservative bounds and enhancing the practicality of robust, safety-and-stability-guaranteeing NN controller design.
I-A Related Work on learning-based control
@@
In imitation learning setting of [2], a human expert remotely controls a drone equipped with a camera to follow a moving object on the ground. The video recording of the human expert collected during this process is subsequently used to train a neural network to imitate the expert’s command for the corresponding raw inputs. In the reinforcement learning (RL) setting of [6], a deterministic on-policy gradient method was applied to train a NN controller for a quadcopter to recover its position under harsh disturbances, demonstrating its stabilizing ability, but without any formal guarantee. Studies in [7], comparing deep RL methods such as Deep Deterministic Policy Gradient (DDPG), Trust Region Policy Optimization (TRPO), and Proximal Policy Optimization (PPO) with traditional PID controllers found that RL-trained NN controllers outperformed their PID counterparts in attitude control tasks. These findings highlight the potential of NN controllers for quadcopter control, often achieving performances better than the conventional methods.
@@
However, a critical limitation of these approaches is their lack of formal guarantee of stability and robustness to environmental and model uncertainties. Such limitation hinder the transferability of NN controllers trained in one environment to another. To address this, researchers have explored methods to robustify the training of NN controllers. Domain randomization, for example, has been employed to account for uncertainties by randomizing model/disturbance parameters during the training. This method enabled a quadcopter to land autonomously at a target point [8], with results showing that the adequately randomized domains allowed the trained policy to generalize to new environments without any additional training. Similarly, the Robust Markov Decision Process (RMDP) was used in [9] to handle environmental and model uncertainties, where RMDP maximizes cumulative rewards under worst-case scenarios. Quadcopter attitude control trained using RMDP demonstrated satisfactory trajectory tracking in uncertain environments [9]. While these methods improve robustness, they lack any formal guarantee.
From a control theory perspective, some studies have analyzed NN controllers to ensure stability. For instance, [10] introduced a quadratic constraint for Lipschitz functions, formulating a semidefinite programming problem to guarantee input-output stability. [1] developed quadratic constraint for Lipschitz-bounded NN controllers, sector bounded nonlinear components of the system and bound for its variability and disturbances, deriving linear matrix inequality (LMI) constraints for robust Lyapunov stability.
@@
To the best of our knowledge, our work is the first that leverages these advancements of robust, safe, and stability-gauranteed NN control design for the navigation of quadcopters for optimal reference trajectory tracking. Our contribution further includes enhancing the method of [1] to allow for less conservative Lipschitz bound by using 2-norm in place of infinity-norm while preserving the key results of robust stability, and by estimating the Lipschitz bound of a NN controller using a less conservative method involving semi-definite programming.
@@ I-B Organization
The remainder of this paper is organized as follows. Section II briefly discusses quadcopter dynamics. Section III reviews the robust optimal safe and stability-guaranteed training (ROSS-GT) method of [1]. Section IV presents the enhancement and application of ROSS-GT for a quadcopter. Section V provides the simulation results of the trained NN controller. Section VI then provides a conclusion.
@@ I-C Notation
For a vector $x\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th element. $\left\lVert x\right\rVert_{p}$ denotes vector’s $p$ norm when $p\geq 1$ and $\left\lVert x\right\rVert_{\infty}$ denotes the infinity norm, $diag(x)$ denotes a diagonal matrix whose $i$-th diagonal element is $x_{i}$. For a matrix $M\in\mathbb{R}^{m\times n}$, $M_{ij}$ denotes its element in $i$-th row and $j$-th column. A symmetric matrix is denoted by $P=\left[\begin{array}[]{cc}P_{11}&*\\ P_{21}&P_{22}\end{array}\right]$. $\mathbb{S},\mathbb{S}^{\geq 0}$, and $\mathbb{S}^{+}$ represent the set of symmetric, positive semi-definite, and positive definite matrices, respectively.
@@ II Quadcopter Model
In this section, quadcopter dynamics is briefly discussed. Quadcopter dynamics consists of translation and rotation movement. In the case of translation movement, gravitation and thrust contribute to the acceleration. Also, the drag force, which is proportional to the velocity, also impacts the translation motion.
@@
where $S_{a}$ and $C_{a}$ indicates $sin(a)$ and $cos(a)$, respectively.
@@
In this work, an X-shape quadcopter is considered. Thus, the thrust and torque values are generated by the individual propeller’s rotations and thrust and drag constants:
@@
III Robust Optimal Safe and Stability Guaranteed Training Method
We apply the Robust Optimal Safe and Stability Guaranteed Training (ROSS-GT) method proposed in [1] to certify the safety and asymptotic stability of neural network (NN)-controlled closed-loop systems with parameter variations and environmental disturbances. Robust asymptotic stability is guaranteed by finding a robust Lyapunov function for the system, which provides a sufficient condition for stability, and a neighborhood of the equilibrium that is a sublevel set of the Lyapunov function serves as a safe invariant domain for initialization.
III-A System and Control Framework
@@
where $s(t)\in\mathbb{R}^{n},u(t)\in\mathbb{R}^{m}$, and $\alpha(t)\in\mathbb{R}^{d}$ are state, control input, and parameter variation, respectively; $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ denotes the nonlinear plant dynamics assumed to be locally continuously differentiable; $\pi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes a state-feedback control policy that is viewed as composition of nominal $\pi_{0}(\cdot)$ and NN $\pi_{NN}(\cdot)$ controllers.
An equilibrium point is defined as the state where the system dynamics remains unchanged ($\dot{s}=0$). The equilibrium point is assumed invariant under parameter variations, and without loss of generality, the coordinates are shifted such that the equilibrium point aligns with the origin ($s^{*}=\boldsymbol{0}$). The system (41) is deemed robustly asymptotically stable if, starting from an initial state within a certain neighborhood of the equilibrium, it converges to the equilibrium point in spite of the parametric variations, assumed to be bounded within a hyper-rectangular, providing upper and lower bounds for the parameters.
When parameter variations are zero, the system is referred to as the nominal system, and the corresponding control is termed the nominal control, and is designed to stabilize the nominal system linearized around the equilibrium point. The NN control is thus viewed as a perturbation around the nominal control to ensure the stability of the overall parameter-varying system (and not just the linearized nominal system). Fig 1 illustrates the structure of the control architecture, which integrates both the nominal and NN controllers. Both the nominal and NN controllers operate as state-feedback controllers, with the plant receiving control commands as the sum of their respective outputs.
Figure 1: Control Architecture
@@
In [1], a feasibility condition for the existence of a Lyapunov function comes from two quadratic constraints (QC) obtained from: 1) NN controller Lipschitz bound, and 2) nonlinear and parameter variation (NPV) sector bound. As demonstrated in [10], Lipschitz continuous functions, such as the NNs, exhibit QCs derived from their bounded partial derivatives within a specified domain. Similarly, sector bounded NPVs imply another set of QCs. By combining these QCs, [1] proposed a Linear Matrix Inequality (LMI) constraint to be able to certify the existence of a Lyapunov function (the details are given below in this section itself).
The ROSS-GT method of [1] is implemented in two stages:
1. Extraction of maximal Lipschitz bound and robust and safe region of stability (RS-RoS): From the given system dynamics with bounds for parameter variations and a safety domain, the maximal RS-RoS, the maximal Lipschitz bound for the NN controller, and a state-feedback gain of nominal controller are derived, iteratively, based on the feasibility of the LMI constraint.
2. Optimal NN Controller Training: A NN controller is then trained using reinforcement learning while ensuring that its Lipschitz constant remains within the bound established in the first stage, and the cost of control and deviation from reference trajectory is minimized.
The original approach in [1] used the infinity norm for defining Lipschitz continuity and deriving QCs associated with the NN control. In this context, the Lipschitz constant of the NN controller was estimated using the maximum absolute row sum norm of its weight matrices. However, this approach often yields overly conservative estimates, particularly for high-dimensional networks, making it impractical. To address this, and make the approach of [1] practical, here we propose replacing the infinity norm with the Euclidean (L2) norm, proving that the feasibiliy of the LMI constraint is preserved under this change. Additionally, we employ a semidefinite programming-based method[5] to obtain tighter Lipschitz estimates, resulting in more practical relaxed Lipschitz bounds for the NN controllers.
III-B Lipschitz Continuity of NN Controller and its QC
A neural network (NN) is a Lipschitz operator, meaning its output magnitude is bounded by the product of the Lipschitz constant and the input magnitude. The Lipschitz continuity of an NN can be defined using any type of norm; in this work, we adopt the L2-norm for the Lipschitz continuity.
Definition 2 (Lipschitz continuity) A function $\pi$ is locally Lipschitz continuous over a domain $\mathcal{S}$ if there exists a constant $L>0$ such that:

### v2

@@ Abstract
Recent advances in deep learning have provided new data-driven ways of controller design to replace the traditional manual synthesis and certification approaches. However, employing neural networks (NN) as controllers presents its own challenge: that of certifying safety, stability and robustness, due to their inherent complex nonlinearity, and while NN controllers have demonstrated high performance in complex systems, they often lack formal guarantees. This issue is further accentuated for critical applications such as unmanned aerial vehicles (UAVs), whereas a lack of safety or stability or robustness assurance raises the risk of critical damage or even complete failure. In this study, we improve a Robust, Optimal, Safe, and Stability Guaranteed Control (ROSS-GC) method of [1] to design an NN controller for a quadcopter flight control. The approach ensures robust closed-loop system stability by finding a robust Lyapunov function and providing a safe initial state domain that remains invariant under the control and also guarantees asymptotic stability to an equilibrium within it even under random parametric variability. Robust stability guaranteeing constraints are derived from the sector bound of the system nonlinearity and of its parametric variations, in the form of a Lipschitz bound for NN control. The control performance is next optimized by searching over the class of controllers having the said Lipschitz bound to minimize the reference tracking error together with the control costs.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I INTRODUCTION
In recent decades, neural networks (NNs) based data-driven approach has achieved great progress, finding applications across diverse fields, including tne promising use case of the design of NN controllers for dynamical systems, where a neural network maps state observations to actions. Unlike conventional feedback controllers, which rely either on predefined structures with domain-specific expertise for their tuning, or the control Lyapunov functions that require their exploration, NN controllers are trained through data-driven approaches. For instance, supervised imitation learning uses images from a human operator and trains a neural network to imitate those actions [2],[3]. Similarly, in reinforcement learning (RL), policy networks are optimized to maximize the expected cumulative rewards from control sequences [4].
@@
While NN controllers have demonstrated good performance in many applications, their deployment in safety-critical systems, such as unmanned aerial vehicles (UAVs), remains limited due to the lack of formal guarantees of safety, stability, and robustness. Safety corresponds to operting in the safe region of the state-space, stability requires convergence to reference, while robustness if the ability to withstand disturbances, sensor noise, or model uncertainties without loss of safety or stability. In real-world environments of uncertainties, robust control is essential for reliable operation, which again requires a formal certification that, in general, is lacking for NN controllers.
@@
This paper enhances and applies the Robust, Optimal, Safe, and Stability-Guaranteed Control (ROSS-GC) method of [1] to design and train an NN controller for a quadcopter, a widely used form of UAV. The ROSS-GC method consists of two stages: First, it extracts the Lyapunov-based constraints to ensure robust safety and stability for a closed-loop nonlinear system with parametric variability, as well as a safe initial state domain that is invariant, and guarantees converge to equilibrium within in spite of the parametric variations, utilizing the sector bounds of system nonlinearities and its parametric variations, and imposing Lipschitz bound for the neural network (so their output magnitude is bounded by the input magnitude and the Lipschitz bound gain). The maximal permissible robust, safe, and stabilizing Lipschitz bound for the NN controller and the maximal initial safe state domain ensuring asymptotic robust stability under bounded parametric uncertainty are derived by iteratively analyzing a linear matrix inequality (LMI) certificate. Second, it finds an optimal NN controller satisfying the Lipschitz bound obtained in the first step to minimize control cost and reference trajectory deviation error using reinforcement learning.
@@
The original ROSS-GC method [1] employs the infinity norm to derive the Lipschitz bound for the NN controller. However, this approach often yields overly conservative estimates for the NN Lipschitz constant, particularly for NNs with many neurons, making it impractical. In this paper, we propose an enhancement, replacing the infinity norm with the Euclidean norm, demonstrating that it still satisfies the same LMI constraints for the existence of a robust Lyapunov function, in spite of reducing the above-mentioned conservativeness. Furthermore, we utilize a semidefinite programming-based method for the estimation of the Lipschitz constant of a neural network [5] that further reduces the conservativeness of the method in [1] (the result is a $>\!\!2400$-fold reduction in conservativeness for the quadcopter example considered), enabling the practicality of robust, safety, and stability guaranteeing NN controller design.
I-A Related works on learning-based control
@@
In the imitation learning setting of [2], a human expert remotely controls a drone equipped with a camera to follow a moving object on the ground. The video recording of the human expert collected during this process is subsequently used to train a neural network to imitate the expert’s command for the corresponding raw inputs. In the reinforcement learning (RL) setting of [6], a deterministic on-policy gradient method was applied to train an NN controller for a quadcopter to recover its position under harsh disturbances, demonstrating its stabilizing ability, but without any formal guarantee. Studies in [7], comparing deep RL methods such as Deep Deterministic Policy Gradient (DDPG), Trust Region Policy Optimization (TRPO), and Proximal Policy Optimization (PPO) with traditional PID controllers have found that RL-trained NN controllers outperform their PID counterparts in attitude control tasks. These findings highlight the potential of NN controllers for quadcopter control.
@@
However, a critical limitation of these approaches is their lack of formal guarantee of safety, stability, and robustness to environmental and model uncertainties. To address this, researchers have explored methods to robustify the training of NN controllers. Domain randomization, for example, has been employed to account for uncertainties by randomizing model/disturbance parameters during the training. This method enabled a quadcopter to land autonomously at a target point, with results showing that the adequately randomized domains allowed the trained policy to generalize to new environments without any additional training [8]. Similarly, the Robust Markov Decision Process (RMDP) was used in [9] to address environmental and model uncertainties, where RMDP maximizes cumulative rewards under worst-case scenarios. Quadcopter attitude control trained using RMDP demonstrated satisfactory trajectory tracking in uncertain environments [9]. While these methods improve robustness, they still lack any formal guarantee.
From a control theory perspective, some studies have analyzed NN controllers to ensure safety, stability, and robustness. A common approach involves employing a Lyapunov function, which provides a sufficient condition for asymptotic stability. One approach is to model the Lyapunov function as well as the controller as neural networks and then train them to certify closed-loop stability [10, 11, 12, 13, 14]. The setup consists of a learner and a verifier, where the verifier validates the Lyapunov constraint and the learner updates the weights of the Lyapunov and the controller NNs. The verifier, using techniques such as mixed-integer programming (MIP) or satisfiability modulo theories (SMT), identifies counterexamples violating stability constraints. If no counterexamples are found, the Lyapunov constraint is considered satisfied. The learner uses counterexamples to tune the weights in the Lyapunov function and the controller. Simulation in [10] showed that the trained controller successfully stabilized a quadcopter in a hovering position. However, such a validation of the Lyapunov constraint is undecidable in general, and the empirical approaches are sound but not complete, and also, the computational complexity of MIP/SMT solvers during the training process poses challenges for adoption to large-scale problems.
Another control-theoretic approach leverages quadratic constraints (QCs) to bound the output of NN controllers, abstracting activation function properties [15, 16, 17, 1]. The method in [16] formulated an LMI constraint to approximate the input-output map of a neural network. When the input domain with disturbance is given, safety and robustness are proved by showing that the output approximation lies within a safe region. To prove robust closed-loop stability with NN controller, [17] derived LMI constraint from integral quadratic constraints (IQCs) that capture the input-output behavior of sector-bounded nonlinear components of the system and QCs that bound the nonlinear activation function in NN controller. However, this method is mainly suited for static networks, and also is difficult to integrate into the training process. An approach in [15] derives QCs from sector bounds of smooth functions to guarantee input-output stability. Extending this idea, [1] developed QCs for Lipschitz-bounded NN controllers, sector-bounded nonlinear components of the system, and its variability, deriving LMI constraints for robust Lyapunov stability. From the proposed LMI, a Lipschitz constraint for an NN controller that can maintain robust stability is obtained. Then, the NN controller is trained while maintaining that its Lipschitz bound lies below the derived constraint.
@@
The methods using QCs represent nonlinear activation functions using constraints that capture an activation function’s property, such as the resulting Lipschitz bound. To obtain a practical solution, however, it is important to obtain non-conservative Lipschitz bound constraints, and we propose certain such enhancements in the paper, along with their application to quadcopter control for the first time. Our work leverages these advancements of robust, safe, and stability-guaranteed NN control design for the navigation of quadcopters for optimal reference trajectory tracking. Our contribution further includes enhancing the method of [1] to allow for a less conservative Lipschitz bound by using 2-norm in place of infinity-norm while preserving the key results of robust stability and by estimating the Lipschitz bound of a NN controller using a less conservative method involving semi-definite programming.
@@ I-B Organization
The remainder of this paper is organized as follows. Section II briefly discusses quadcopter dynamics. Section III reviews the robust optimal safe and stability-guaranteed control (ROSS-GC) method of [1]. Section IV presents the enhancement and application of ROSS-GC for a quadcopter. Section V provides the simulation results of the trained NN controller. Section VI then provides a conclusion.
@@ I-C Notation
For a vector $x\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th element. $\left\lVert x\right\rVert_{p}$ denotes vector’s $p$ norm when $p\geq 1$ and $\left\lVert x\right\rVert_{\infty}$ denotes the infinity norm, $diag(x)$ denotes a diagonal matrix whose $i$-th diagonal element is $x_{i}$. For a matrix $M\in\mathbb{R}^{m\times n}$, $M_{ij}$ denotes its element in $i$-th row and $j$-th column. Operators ,<$\leq$,<$\geq$ , < , ≥, and $>$ on matrices or vectors are elementwise operations. For simplicity of notation, a symmetric matrix is denoted by $P=\left[\begin{array}[]{cc}P_{11}&*\\ P_{21}&P_{22}\end{array}\right]$. $\mathbb{S},\mathbb{S}^{\geq 0}$, and $\mathbb{S}^{+}$ represent the set of symmetric, positive semi-definite, and positive definite matrices, respectively.
@@ II Quadcopter Model
In this section, quadcopter dynamics is briefly discussed. Quadcopter dynamics consists of translation and rotation movement. In the case of translation movement, gravitation, thrust, and drag force, which is proportional to the velocity, impact the translational acceleration.
@@
in which $S_{a}$ and $C_{a}$ indicates $sin(a)$ and $cos(a)$, respectively.
@@
The angular velocity in body frame and Euler angle rate are related by the equality (37):
$\left[\begin{array}[]{c}\dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\end{array}\right]=\left[\begin{array}[]{ccc}1&S_{\phi}T_{\theta}&C_{\phi}T_{\theta}\\ 0&C_{\phi}&-S_{\theta}\\ 0&\frac{S_{\phi}}{C_{\theta}}&\frac{C_{\phi}}{C_{\theta}}\end{array}\right]\left[\begin{array}[]{c}p\\ q\\ r\end{array}\right].$ (37)
In this work, an X-shape quadcopter is considered, for which the thrust and torque values as generated by the individual propeller’s rotations and the drag constants are given by,
@@
III Robust Optimal Safe and Stability Guaranteed Control Method
We review the Robust Optimal Safe and Stability Guaranteed Control (ROSS-GC) method proposed in [1] to certify the robust safety and asymptotic stability of neural network (NN)-controlled closed-loop systems with random parameter variations.
III-A System and its control framework
@@
where $s(t)\in\mathbb{R}^{n},u(t)\in\mathbb{R}^{m}$, and $\alpha(t)\in\mathbb{R}^{d}$ are state, control input, and parameter variation, respectively; $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ denotes the nonlinear plant dynamics assumed to be locally continuously differentiable; $\pi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes a state-feedback control policy that is viewed as a composition of a nominal $\pi_{0}(\cdot)$ and an NN $\pi_{NN}(\cdot)$ control.
An equilibrium point is defined as the state where the system dynamics remain unchanged ($\dot{s}=0$). In this study, the equilibrium point is assumed invariant under parameter variations, and without loss of generality, the coordinates are shifted such that the equilibrium point aligns with the origin ($s^{*}=\boldsymbol{0}$). The system (51) is deemed robustly asymptotically stable if, starting from an initial state within a certain neighborhood of the equilibrium, it converges to the equilibrium point in spite of the parametric variations, assumed to be contained within a bounded domain.
Figure 1: Controller with nominal vs. NN components
When parameter variations are zero, the system is referred to as the nominal system, and the corresponding control is termed the nominal control, and is designed to stabilize the nominal system linearized around the equilibrium point. The NN control is thus viewed as a perturbation around the nominal control to ensure the stability of the overall nonlinear parameter varying system (and not just the linearized nominal system). Fig 1 illustrates this structure of the control architecture, which integrates both the nominal and NN controllers. Both the nominal and NN controllers operate as state-feedback controllers, with the plant receiving control command as the sum of their respective outputs.
@@
In [1], a feasibility condition for the existence of a Lyapunov function comes from two quadratic constraints (QC) obtained from: 1) NN controller Lipschitz bound, and 2) system’s nonlinear and parameter variation (NPV) sector bound. As demonstrated in [15], Lipschitz continuous functions, such as the NNs, exhibit QCs derived from their bounded partial derivatives within a specified domain. Similarly, sector bounded NPVs imply another set of QCs. By combining these QCs, [1] proposed a Linear Matrix Inequality (LMI) constraint to be able to certify the existence of a robust Lyapunov function, as discussed in the following.
The ROSS-GC method of [1] consists of two stages:
1. Extraction of maximal Lipschitz bound, robust and safe region of stability, and nominal control: From the given system dynamics with bounds for parameter variations and a safety domain of operation, an LMI constraint for the existence of robust Lyapunov function is developed, and using which, the maximal RS-RoS, the maximal Lipschitz bound for the NN controller, and a state-feedback gain of a nominal controller are derived, based on the iterative feasibility check of the LMI constraint.
2. Optimal NN Controller Training: A NN controller is then trained using reinforcement learning while ensuring that its Lipschitz constant remains within the bound established in the first stage, and the cost of control, as well as the deviation of the state trajectory from a desired reference, is minimized.
The approach in [1] used the infinity norm for defining Lipschitz continuity and deriving QCs associated with the NN control. Further, the Lipschitz constant of the NN controller was estimated using the maximum absolute row sum norm of its weight matrices. However, this approach often yields an overly conservative estimate for the Lipschitz bound, particularly for high-dimensional networks, making it impractical. To address this, to be able to make the approach of [1] practical, here we propose replacing the infinity norm with the Euclidean (L2) norm, proving that the feasibility of the LMI constraint remains preserved under this change. Additionally, we employ a semidefinite programming-based method[5] to obtain a tighter Lipschitz bound estimate, resulting in a more practical, relaxed Lipschitz bound for the NN controllers.
III-B Lipschitz bounded control with New enhancements
A neural network (NN) that we use for control is a Lipschitz operator, meaning its output magnitude is bounded by the product of the input magnitude and a certain constant that depends on the NN parameters. This Lipschitz continuity of an NN can be defined using any norm; in this work, we adopt the L2-norm for the Lipschitz continuity, replacing the $\infty$-norm used in [1]. This change preserves all the results of [1], yet makes the Lipschitz bound much less conservative, and so far more easier to satisfy.
Definition 2 (Lipschitz continuity) A function $\pi$ is Lipschitz continuous over a domain $\mathcal{S}$ if there exists a constant $L>0$ such that: